# Astrophysical Implications of Eccentricity in Gravitational Waves from Neutron Star-Black Hole Binaries

### v1

@@ II Methods
For both detectability and measurability studies presented here, we use the gravitational waveform model pyEFPE [pyEFPE], an inspiral-only approximant containing the effects of eccentricity, a variable mean anomaly parameter, and misaligned spins. Studies with pyEFPE were the first to demonstrate evidence for eccentricity in NSBH signal GW200105 [2025arXiv250315393M]. pyEFPE is computationally efficient, making it appropriate for population recovery studies requiring $\mathcal{O}(10^{4})$ or more waveforms to be generated. While pyEFPE is limited by being an inspiral-only waveform model without HMs, merging NSBHs have inspiral-dominated signals, and confident NSBH signals so far have not demonstrated evidence for HMs. To assess the impact of including higher-order modes on the measurability of eccentricity, we use SEOBNRv5EHM [SEOBNRv5EHM, SEOBNRv5EHM:theory], an effective one-body full inspiral-merger-ringdown waveform model with eccentricity and aligned spins, that has also been used to demonstrate evidence for eccentricity in GW200105 [Kacanja:2025:NSeventseccentric, 2025arXiv250812460J].
@@
We generate waveforms with minimum and reference frequencies of $10$ Hz, and inject them into simulated O4-sensitivity LVK detector data using the Bayesian inference library Bilby [bilby, bilby2]. When generating waveforms, we pass in eccentricity defined at an orbit-averaged frequency of $f_{22}=10$ Hz.
II.1 Measurability of eccentricity in NSBHs
@@
We follow Lower2018 to estimate the threshold at which eccentricity becomes detectable:
@@
II.2 Detectability of eccentric NSBHs
Figure 1: Left: The translucent blue filled histogram shows the eccentricity defined at a peak GW frequency of $10$ Hz, output by the triple simulations of StegmannKlencki2025. The translucent orange filled histogram shows these same eccentricities but evolved to the $10$ Hz $22$-mode frequency using the standard eccentricity definition scripts of Vijaykumar2024 (see also 2023:Shaikh:ecc) or the analytic solutions of 2021PhRvD.104j4023T, assuming no redshifting. The green unfilled histogram shows the source-frame eccentricity of systems that are recovered by the optimal quasi-circular aligned-spin search described in the text, averaging every injected system over $500$ possible $d_{L}\ell$ pairs and $10^{7}$ sets of extrinsic parameters, taking into account the effects of redshifting in the injection. Our mock quasi-circular search recovers only $27\%$ of the injected signals. The vertical purple line shows the median $e_{22,10}$ recovered in 2025arXiv250315393M for GW200105 using the same waveform model as is used here, and the translucent purple band spans the $90\%$ credible interval around this median. Right: The found percentage of events with $\log_{10}(e^{\prime}_{10})$ equal to or greater than the value shown on the horizontal axis.
@@ II.2 Detectability of eccentric NSBHs
We take the parameters of the merging NSBHs from the massive triple star simulations of StegmannKlencki2025; interested readers should consult that paper for full details of the simulations. The $N_{\rm sim}=1542$ NSBH mergers in these simulations are representative results from the evolution of a wide range of isolated triple star systems, because only a narrow subset of such systems produces a merging NSBH binary.
@@
Each simulated system comprises three stars: two in an inner binary with masses $m_{*,1}\geq 10$ M, $m_{*,2}\geq 5$ M, and one less massive tertiary with $m_{*,3}<5$ M. The mass of the primary is sampled from a Kroupa2001 mass function in the range $10\leq m_{1}\leq 100$ M. There is no correlation between the black hole mass $m_{1}$ and the final eccentricity at $10$ Hz. When the secondary becomes an NS, its mass is fixed to $m_{2}=1.5$ M. The simulations are carried out at a fixed metallicity $Z\approx 0.1Z_{\odot}$: for isolated NSBH mergers, metallicity has a weak impact on merger rates and mass distributions [e.g., 2019MNRAS.490.3740N, 2021ApJ...912L..23R, 2022MNRAS.516.5737B, 2023MNRAS.524..426I], and it is assumed here that the same is true for triples.
@@
The inner binary evolution proceeds in a two-stage process typical of NSBH formation [2021ApJ...920L..13B, 2024A&A...683A.144X]: (i) the primary becomes a BH while the secondary is still on the main sequence, then (ii) the OB secondary becomes an NS. The birth of the NS—the most disruptive step in the evolution of these systems—acts to filter out triples that are not closely bound. There are two main causes of disruption: first, the inner binary could be disrupted due to the natal kick of the newborn NS; second, the recoil kick to the centre-of-mass of the nascent NSBH due to the mass loss of the secondary could detach the binary from the tertiary, with a kick magnitude roughly inversely proportional to the inner binary separation $a_{1}$ [cf. LuNaoz2019]. StegmannKlencki2025 find that the natal kick imposes an strict upper limit on the inner orbital separation ($a_{1}\lesssim 10^{2}\,\rm AU$), while only weakly restricting the outer orbit ($a_{2}\lesssim 10^{4}\,\rm AU$) as a result of the unequal mass ratio of the newly born NSBH. However, the distribution of surviving NSBH triples is strongly skewed by the requirements imposed by the recoil kick to the centre-of-mass: the tighter the inner binary, the closer the tertiary must be in order for the system to remain bound as a triple. For the surviving systems, $a_{2}/a_{1}$, peaks at $\approx 10$ with a sharp cut-off at $10^{3}$.
@@
As a result, the surviving NSBHs are confined in closely-bound, compact triples. The strength of the octupole term in the secular equations of motion for hierarchical triples scales with both $a_{1}/a_{2}$ and (1q)/(1+q)(1-q)/(1+q) [2013MNRAS.431.2155N]. Consequently, due to the generally more unequal mass ratios of NSBHs, octupole terms in the evolution become more important [Liu2015, Naoz2016, StegmannKlencki2025]. This makes the occurrence of ZKL oscillations less conditional on the outer binary plane being close-to-perpendicular to the tertiary orbital plane, thereby giving rise to higher eccentricities from a wide range of initial conditions [StegmannKlencki2025].
@@
Using scripts from Vijaykumar2024, we convert the simulation-output eccentricity defined at the Wen2003 peak GW frequency of $10$ Hz, $e_{p,10}$, to an orbit-averaged 22-mode ($l=m=2$) frequency [2023:Shaikh:ecc] of $10$ Hz, $e_{22,10}$. The former frequency definition corresponds to that of the harmonic at which the maximum power is emitted for an eccentric binary, and can be thought of as the GW frequency at periastron [e.g., Wen2003, Hamers2021, Vijaykumar2024]. It is recommended that the latter definition is used instead for GW observations: this increases monotonically over the course of the inspiral, avoids the low- and high-eccentricity breakdowns of the Wen2003 definition, and enables eccentricity to be measured directly from the gravitational waveform [2023:Shaikh:ecc].
@@
Of the 1542 binaries, 70 ($4.5\%$ of the total) cause the conversion script to fail due to numerical issues in those scripts, which occur for systems with very small semi-latus rectum $p=a(1-e^{2})$. For most of these (53 of 70), we are able to use the formulae of 2021PhRvD.104j4023T to evolve to the converted 22-mode eccentricity; for the remaining 17, we discard the systems. The difference between the two definitions can be seen between the blue and orange histograms in the left-hand panel of Figure 1. Discarded systems are included in the histogram of $e_{p,10}$, but not in the histogram of $e_{22,10}$.
@@
We first wish to represent a realistically-distributed population of triple-origin NSBH mergers. For each of the $N_{\mathrm{sim}}$ binaries in the simulation output, we generate $N_{\mathrm{pop}}=500$ pairs of luminosity distance $d_{\mathrm{L}}$ and relativistic anomaly $l$ values. We assume a uniform distribution between 0 and $2\pi$ for $l$, and a uniform in source-frame distribution from $1$ to $1000$ Mpc for $d_{\mathrm{L}}$. [^1] The maximum distance to which the heaviest system output by the simulation can be detected is just below $1000$ Mpc; at $1000$ Mpc, this system has a matched-filter SNR of $\simeq 6$ with $e_{22,10}=0$.
@@
We then build on the method used in Zevin2021 to compute the detectability of our simulated population.
@@
For each $i$ of the $N_{\mathrm{pop}}$ pairs, we inject the eccentric NSBH signal with the optimal source orientation of face-on, directly overhead, and perpendicularly polarised. We compute the maximum matched-filter signal-to-noise ratio (SNR), $\rho_{\mathrm{max},i}$, obtained with the quasi-circular waveform template. We set the sampling frequency to $4096$ Hz, and allow $128$ s of data when generating waveforms.
@@
While Zevin2021 computed the SNR against a quasi-circular waveform with all other parameters matching, the SNRs of inspiral-dominated NSBH signals are far more sensitive to small changes between the matched templates. Before computing the SNR, we therefore minimise the mismatch between the eccentric signal and quasi-circular template by varying the geocent time $t_{0}$, phase $\phi$, polarisation $\psi$, and amplitude $\mathcal{A}$. We also set the $x$ and $y$ components of the primary’s spin vector to 0 in the quasi-circular template. Other parameters of the quasi-circular aligned-spin template waveform match those of the eccentric and arbitrarily spinning injection. Attempts to minimise the mismatch over mass and spin parameters yielded negligible improvements over just optimising $(t_{0},\phi,\psi,\mathcal{A})$, so were left out of final results. Our maximum matched-filter SNRs may still be lower than the true possible maximum, since we do not optimise over the whole range of intrinsic parameters. The detection efficiency we estimate here should therefore be taken as a lower limit.
@@
We then draw $N_{\mathrm{ext}}=10^{7}$ realisations of extrinsic parameters other than $d_{\mathrm{L}}$ and $l$ from the default bilby prior distributions for binary BHs [bilby2]. We compute the detector projection factor $\Theta$ for each $j$ of these realisations 1993PhRvD..47.2198F:
@@
$P_{\mathrm{det},i}=\sum_{j=1}^{N_{\mathrm{ext}}}\mathcal{H}\left[\frac{\Theta_{j}}{4}\rho_{\mathrm{max}}-\rho_{\mathrm{thresh}}\right],$ (4)
@@
Note that we assess the detectability of signals from NSBH sources based only on their matched-filter SNR, which is the optimal ranking statistic in Gaussian noise. In reality, templated searches must contend with non-Gaussian transients (“glitches”) that can produce spuriously high SNRs. To mitigate the impact of this, searches evaluate the consistency of the data with the template using methods that look beyond the raw SNR; for example, by checking whether the SNR accumulates with frequency as predicted by the waveform model [e.g., Allen:2004gu, 2012PhRvD..85l2006A]. Since the frequency evolution of eccentric signals deviate from quasi-circular templates, these consistency checks tend to down-weight their triggers even if the SNR is high. Because our analysis does not account for this signal–template inconsistency penalty, our detectability estimates for a triple-origin eccentric NSBH population may be over-optimistic.
III Results
III.1 Minimum measurable eccentricities for NSBHs
Figure 2: Scatter plot of total mass vs estimated minimum measurable eccentricity calculated via the method described in the text, for injections of pyEFPE waveforms with either minimum or maximum values of $\chi_{\mathrm{p}}$. The H1 SNR of the quasi-circular versions of these injections is within $\pm 0.4$ of $\rho_{0,H1}=22$. For a fixed SNR, the minimum detectable eccentricity correlates with total mass when $\chi_{\mathrm{p}}=0$. When $\chi_{\mathrm{p}}$ is maximal ($\chi_{1}^{\perp}=0.99$; we assume the secondary is non-spinning), the relationship is not linear, and measurability improves for more extreme mass ratios. The colour range indicates the mass ratio $q$.
@@ III.1 Minimum measurable eccentricities for NSBHs
III.2 Astrophysical implications of detecting eccentric NSBHs
@@
In the left-hand panel of Figure 1, we compare the histograms of the simulation-output source-frame eccentricities with a histogram of the source eccentricities of injections found using our quasi-circular aligned-spin mock search, including the effects of redshifting and marginalising over variations in extrinsic parameters. Our mock search recovers $27\%$ of the total simulated population. If instead we compute the mock search using eccentric and spin-precessing templates, we recover $52\%$ of our total simulated population. Alternatively, if we alter our simulated population to have only aligned spins and zero eccentricity, the aligned-spin quasi-circular mock search recovers $64\%$ of systems. The relative detection efficiency of the search is therefore $52\%$ compared to a search containing the same physics as the injected population, and $42\%$ compared to the recovery of an equivalent non-eccentric non-spin-precessing population.
@@
We find that the search sensitivity drops steeply above $e_{22,10}=0.01$, and that in the region of the eccentricity posterior for GW200105, there are about three times as many missed signals as found signals. In the right-hand panel of Figure 1, we show the percentage of signals that are found with an eccentricity greater than or equal to $e_{22,10}$. When $e_{22,10}\geq 0.01$, only $13\%$ of sources are recovered.
@@
In Zevin2021, $e_{10}=0.05$ was used as the minimum detectable eccentricity threshold for binary BH mergers, based on the recovery of GW150914-like injections using Bayesian inference [Lower2018]. Using the optimistic overlap method, we estimate that NSBH-like systems have similar detectable eccentricities of $e_{22,10}\approx\mathcal{O}(0.01)$. In fact, the overlap estimates of Lower2018 for GW150914-like systems are slightly lower than our estimates for NSBHs, possibly due to using a different waveform model or the higher SNR of the GW150914-like injections.
@@
Fortunately, we do not require an accurate and universal “minimum detectable eccentricity” to establish the astrophysical implications of detecting eccentric NSBHs; we merely require an arbitrary eccentricity threshold above which we are confident that eccentricity measurement would be achieved. In accordance with standard practice for binary BH eccentricity measurability, we choose $e_{22,10}=0.05$ [Lower2018, Zevin2021]. In the simulations of StegmannKlencki2025, most of the NSBH mergers have eccentricities above this threshold: $\approx 80\%$ of NSBHs have $e_{22,10}\gtrsim 0.05$. Due to the low detectability of highly-eccentric inspiral-dominated systems, though, only $19\%$ of these are detected in our mock search. We therefore set the “above threshold” eccentric fraction, i.e. the fraction of systems that we are confident are detectable and measurably eccentric, to $\xi_{\mathrm{thresh}}=0.80\times 0.19=0.152$. Using the same approach and notation as Zevin2021, in Figure 3, we plot the symmetric $95\%$ credible intervals around $\beta_{t}$, the branching fraction of field triples to the total population of observed NSBHs. If the number of observed NSBHs is $N_{\mathrm{obs}}=3$, and if GW200105 is truly eccentric so $N_{\mathrm{ecc}}=1$, then field triples may constitute anywhere between $5\%$ and $100\%$ of observed NSBHs, which reflects a large uncertainty due to the current low-number statistics.
@@
There are likely more than three NSBHs in the catalogue so far. In addition to GW190814, which may be an NSBH with an unusually massive NS secondary, GWTC-2 included the low-significance NSBH candidate GW190426_152155 [GWTC-2]. In GWTC-2.1, one confident candidate, GW190917_114630, has a secondary consistent with being an NS ($m_{2}\approx 2.1$ M), although it is identified by the search pipeline as a binary BH; there are two further low-significance NSBH candidates, GW190531_023648 and GW190917_114630 [GWTC-2-1]. In GWTC-3, in addition to GW200115 and GW200105, two potential NSBHs—GW191219_163120 and GW200210_092254—were reported, although the former has an uncertain astrophysical probability, and the latter, like GW190814, has an ambiguous secondary mass [GWTC-3]. In GWTC-4, both GW230529 and confident NSBH candidate GW230518_125908 were reported [GWTC-4]; however, the latter was not included in downstream analyses (e.g., the rates and populations analysis [GWTC-4-astro-pop]) due to being detected outside of official observing, during the engineering run.
@@
We include in Figure 3 a vertical red band spanning $3\leq N_{\mathrm{obs}}\leq 11$, demonstrating the uncertainty on the number of NSBH detections. With $N_{\mathrm{obs}}=11$ and $N_{\mathrm{ecc}}=1$, the lower $90\%$ credible limit on $\beta_{t}$ is pushed down to $\approx 1\%$, while the detections remain consistent with $100\%$ of observed NSBHs originating in field triples at $90\%$ credibility. If no more eccentric NSBHs are detected, the upper $90\%$ credible limit on $\beta_{t}$ can begin to be constrained to <1<1 when we reach $N_{\mathrm{obs}}=24$.

### v2

@@ 2 Methods
For both detectability and measurability studies presented here, we use the gravitational waveform model pyEFPE (Morras et al., 2025a), an inspiral-only approximant containing the effects of eccentricity, a variable mean anomaly parameter, and misaligned spins. Studies with pyEFPE were the first to demonstrate evidence for eccentricity in NSBH signal GW200105 (Morras et al., 2025b). pyEFPE is computationally efficient, making it appropriate for population recovery studies requiring $\mathcal{O}(10^{4})$ or more waveforms to be generated. While pyEFPE is limited by being an inspiral-only waveform model without HMs, merging NSBHs have inspiral-dominated signals, and confident NSBH signals so far have not demonstrated evidence for HMs. To assess the impact of including higher-order modes on the measurability of eccentricity, we use SEOBNRv5EHM (Gamboa et al., 2025a, b), an effective one-body full inspiral-merger-ringdown waveform model with eccentricity and aligned spins, that has also been used to demonstrate evidence for eccentricity in GW200105 (Kacanja et al., 2025a; Jan et al., 2025).
@@
We generate waveforms with minimum and reference frequencies of $10$ Hz, and inject them into simulated O4-sensitivity LVK detector data using the Bayesian inference library Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020a). When generating waveforms, we pass in eccentricity defined at an orbit-averaged frequency of $f_{22}=10$ Hz.
2.1 Measurability of eccentricity in NSBHs
@@
We follow Lower et al. (2018) to estimate the threshold at which eccentricity becomes detectable:
@@
2.2 Detectability of eccentric NSBHs
Figure 1: Left: The translucent blue filled histogram shows the eccentricity defined at a peak GW frequency of $10$ Hz, output by the triple simulations of Stegmann and Klencki (2025). The translucent orange filled histogram shows these same eccentricities but evolved to the $10$ Hz $22$-mode frequency using the standard eccentricity definition scripts of Vijaykumar et al. (2024) (see also Shaikh et al. (2023)) or the analytic solutions of Tucker and Will (2021), assuming no redshifting. The green unfilled histogram shows the source-frame eccentricity of systems that are recovered by the optimal quasi-circular aligned-spin search described in the text, averaging every injected system over $500$ possible $d_{L}\ell$ pairs and $10^{7}$ sets of extrinsic parameters, taking into account the effects of redshifting in the injection. The grey histogram shows the results of the same search, but this time including both eccentricity and spin-precession (E+P) in the template waveforms. Our mock quasi-circular search recovers $3.9\%$ of the simulated population, representing $46\%$ of signals recovered in the “E+P” search, which recovers $8.5\%$. The vertical purple line shows the median $e_{22,10}$ recovered in Morras et al. (2025b) for GW200105 using the same waveform model as is used here, and the translucent purple band spans the $90\%$ credible interval around this median. Right: The found percentage of events with $\log_{10}(e^{\prime}_{10})$ equal to or greater than the value shown on the horizontal axis in the quasi-circular search (green) and eccentric spin-precessing search (grey).
@@ 2.2 Detectability of eccentric NSBHs
We take the parameters of the merging NSBHs from the massive triple star simulations of Stegmann and Klencki (2025); interested readers should consult that paper for full details of the simulations. The $N_{\rm sim}=1542$ NSBH mergers in these simulations are representative results from the evolution of a wide range of isolated triple star systems, because only a narrow subset of such systems produces a merging NSBH binary.
@@
Each simulated system comprises three stars: two in an inner binary with masses $m_{*,1}\geq 10$ M, $m_{*,2}\geq 5$ M, and one less massive tertiary with $m_{*,3}<5$ M. The mass of the primary is sampled from a Kroupa (2001) mass function in the range $10\leq m_{1}\leq 100$ M. There is no correlation between the black hole mass $m_{1}$ and the final eccentricity at $10$ Hz. When the secondary becomes an NS, its mass is fixed to $m_{2}=1.5$ M. The simulations are carried out at a fixed metallicity $Z\approx 0.1Z_{\odot}$: for isolated NSBH mergers, metallicity has a weak impact on merger rates and mass distributions (e.g., Neijssel et al., 2019; Román-Garza et al., 2021; Broekgaarden et al., 2022; Iorio et al., 2023), and it is assumed here that the same is true for triples.
@@
The inner binary evolution proceeds in a two-stage process typical of NSBH formation (Broekgaarden and Berger, 2021; Xing et al., 2024): (i) the primary becomes a BH while the secondary is still on the main sequence, then (ii) the OB secondary becomes an NS. The birth of the NS—the most disruptive step in the evolution of these systems—acts to filter out triples that are not closely bound. There are two main causes of disruption: first, the inner binary could be disrupted due to the natal kick of the newborn NS; second, the recoil kick to the centre-of-mass of the nascent NSBH due to the mass loss of the secondary could detach the binary from the tertiary, with a kick magnitude roughly inversely proportional to the inner binary separation $a_{1}$ (cf. Lu and Naoz, 2019). Stegmann and Klencki (2025) find that the natal kick imposes an strict upper limit on the inner orbital separation ($a_{1}\lesssim 10^{2}\,\rm AU$), while only weakly restricting the outer orbit ($a_{2}\lesssim 10^{4}\,\rm AU$) as a result of the unequal mass ratio of the newly born NSBH. However, the distribution of surviving NSBH triples is strongly skewed by the requirements imposed by the recoil kick to the centre-of-mass: the tighter the inner binary, the closer the tertiary must be in order for the system to remain bound as a triple. For the surviving systems, $a_{2}/a_{1}$, peaks at $\approx 10$ with a sharp cut-off at $10^{3}$.
@@
As a result, the surviving NSBHs are confined in closely-bound, compact triples. The strength of the octupole term in the secular equations of motion for hierarchical triples scales with both $a_{1}/a_{2}$ and (1q)/(1+q)(1-q)/(1+q) (Naoz et al., 2013). Consequently, due to the generally more unequal mass ratios of NSBHs, octupole terms in the evolution become more important (Liu et al., 2015; Naoz, 2016; Stegmann and Klencki, 2025). This makes the occurrence of ZKL oscillations less conditional on the outer binary plane being close-to-perpendicular to the tertiary orbital plane, thereby giving rise to higher eccentricities from a wide range of initial conditions (Stegmann and Klencki, 2025).
@@
Using scripts from Vijaykumar et al. (2024), we convert the simulation-output eccentricity defined at the Wen (2003) peak GW frequency of $10$ Hz, $e_{p,10}$, to an orbit-averaged 22-mode ($l=m=2$) frequency (Shaikh et al., 2023) of $10$ Hz, $e_{22,10}$. The former frequency definition corresponds to that of the harmonic at which the maximum power is emitted for an eccentric binary, and can be thought of as the GW frequency at periastron (e.g., Wen, 2003; Hamers, 2021; Vijaykumar et al., 2024). It is recommended that the latter definition is used instead for GW observations: this increases monotonically over the course of the inspiral, avoids the low- and high-eccentricity breakdowns of the Wen (2003) definition, and enables eccentricity to be measured directly from the gravitational waveform (Shaikh et al., 2023).
@@
Of the 1542 binaries, 70 ($4.5\%$ of the total) cause the conversion script to fail due to numerical issues in those scripts, which occur for systems with very small semi-latus rectum $p=a(1-e^{2})$. For most of these (53 of 70), we are able to use the formulae of Tucker and Will (2021) to evolve to the converted 22-mode eccentricity; for the remaining 17, we discard the systems. The difference between the two definitions can be seen between the blue and orange histograms in the left-hand panel of Figure 1. Discarded systems are included in the histogram of $e_{p,10}$, but not in the histogram of $e_{22,10}$.
@@
We first wish to represent a realistically-distributed population of triple-origin NSBH mergers. For each of the $N_{\mathrm{sim}}$ binaries in the simulation output, we generate $N_{\mathrm{pop}}=500$ pairs of luminosity distance $d_{\mathrm{L}}$ and relativistic anomaly $l$ values. We assume a uniform distribution between 0 and $2\pi$ for $l$. We use the UniformSourceFrame prior from bilby to generate our $d_{L}$ distribution in the range $1$ to $1000$ Mpc; this corresponds to a redshift distribution that is uniform in comoving volume and source-frame time. We set this relatively low upper limit on $d_{L}$ to mitigate computational costs; the maximum distance to which the heaviest system output by the simulation can be detected, if we assume $e_{22,10}=0$, is just below $1000$ Mpc (at $1000$ Mpc, this system has a matched-filter SNR of $\simeq 6$).
@@
We then build on the method used in Zevin et al. (2021), which is itself built on the prescriptions in Finn and Chernoff (1993), to compute the detectability of our simulated population.
@@
To start, we calculate the maximum SNR possible for each system, given quasi-circular aligned-spin recovery templates; we will temper this with realistic extrinsic parameter distributions later. For each $i$ of the $N_{\mathrm{pop}}$ pairs, we inject the eccentric NSBH signal with the optimal source orientation of face-on, directly overhead, and circularly polarised. We compute the maximum matched-filter signal-to-noise ratio (SNR), $\rho_{\mathrm{max},i}$, obtained with the quasi-circular waveform template. We set the sampling frequency to $4096$ Hz, and allow $128$ s of data when generating waveforms.
@@
While Zevin et al. (2021) computed the SNR against a quasi-circular waveform with all other parameters matching, the SNRs of inspiral-dominated NSBH signals are far more sensitive to small changes between the matched templates. Before computing the maximum possible SNR, we therefore minimise the mismatch between the eccentric signal and quasi-circular template by varying the geocent time $t_{0}$, phase $\phi$, polarisation $\psi$, and amplitude $\mathcal{A}$ following the method described in Harry et al. (2016). We also set the $x$ and $y$ components of the primary’s spin vector to 0 in the quasi-circular template. Other parameters of the quasi-circular aligned-spin template waveform match those of the eccentric and arbitrarily spinning injection. Attempts to minimise the mismatch over mass and spin parameters yielded negligible improvements over just optimising $(t_{0},\phi,\psi,\mathcal{A})$, so were left out of final results. Our maximum matched-filter SNRs may still be lower than the true possible maximum, since we do not optimise over the whole range of intrinsic parameters. The detection efficiency we estimate here should therefore be taken as a lower limit.
@@
To continue with more realistic SNR estimates, we then draw $N_{\mathrm{ext}}=10^{7}$ realisations of extrinsic parameters other than $d_{\mathrm{L}}$ and $l$ from the default bilby prior distributions for binary BHs (Romero-Shaw et al., 2020a). We compute the detector projection factor $\Theta$ for each $j$ of these realisations (Finn and Chernoff, 1993):
@@
$P_{\mathrm{det},i}=\sum_{j=1}^{N_{\mathrm{ext}}}\mathcal{H}\left[\frac{\Theta_{j}}{4}\rho_{\mathrm{max},i}-\rho_{\mathrm{thresh}}\right],$ (4)
@@
Note that we assess the detectability of signals from NSBH sources based only on their matched-filter SNR, which is the optimal ranking statistic in Gaussian noise. In reality, templated searches must contend with non-Gaussian transients (“glitches”) that can produce spuriously high SNRs. To mitigate the impact of this, searches evaluate the consistency of the data with the template using methods that look beyond the raw SNR; for example, by checking whether the SNR accumulates with frequency as predicted by the waveform model (e.g., Allen, 2005; Allen et al., 2012). Since the frequency evolution of eccentric signals deviate from quasi-circular templates, these consistency checks tend to down-weight their triggers even if the SNR is high. Because our analysis does not account for this signal–template inconsistency penalty, our detectability estimates for a triple-origin eccentric NSBH population may be over-optimistic.
3 Results
3.1 Minimum measurable eccentricities for NSBHs
Figure 2: Scatter plot of total mass vs estimated minimum measurable eccentricity calculated via the method described in the text, for injections of pyEFPE waveforms with either minimum or maximum values of $\chi_{\mathrm{p}}$ (Schmidt et al., 2015). The H1 SNR of the quasi-circular versions of these injections is within $\pm 0.4$ of $\rho_{0,H1}=22$. For a fixed SNR, the minimum detectable eccentricity correlates with total mass when $\chi_{\mathrm{p}}=0$. When $\chi_{\mathrm{p}}$ is maximal ($\chi_{1}^{\perp}=0.99$; we assume the secondary is non-spinning), the relationship is not linear, and measurability improves for more extreme mass ratios. The colour range indicates the mass ratio $q$.
@@ 3.1 Minimum measurable eccentricities for NSBHs
3.2 Astrophysical implications of detecting eccentric NSBHs
@@
In the left-hand panel of Figure 1, we compare the histograms of the simulation-output source-frame eccentricities with a histogram of the source eccentricities of injections found using our quasi-circular aligned-spin mock search, including the effects of redshifting and marginalising over variations in extrinsic parameters. Our mock search recovers $3.9\%$ of the total simulated population. If we compute the mock search using eccentric and spin-precessing templates, we recover $8.5\%$ of our total simulated population. Alternatively, if we alter our simulated population to have only aligned spins and zero eccentricity, the aligned-spin quasi-circular mock search recovers $10.5\%$ of systems. The relative detection efficiency of the search is therefore $46\%$ compared to a search containing the same physics as the injected population, and $37\%$ compared to the recovery of an equivalent non-eccentric non-spin-precessing population.
@@
We find that the search sensitivity drops steeply above $e_{22,10}=0.05$, and that in the region of the eccentricity posterior for GW200105, there are about eight times as many signals that are found in the eccentric and spin-precessing search as found in the quasi-circular search. In the right-hand panel of Figure 1, we show the percentage of signals that are found with an eccentricity greater than or equal to $e_{22,10}$. When $e_{22,10}\geq 0.01$, only $2\%$ of sources are recovered.
@@
In Zevin et al. (2021), $e_{10}=0.05$ was used as the minimum detectable eccentricity threshold for binary BH mergers, based on the recovery of GW150914-like injections using Bayesian inference (Lower et al., 2018). Using the optimistic overlap method, we estimate that NSBH-like systems have similar detectable eccentricities of $e_{22,10}\approx\mathcal{O}(0.01)$. In fact, the overlap estimates of Lower et al. (2018) for GW150914-like systems are slightly lower than our estimates for NSBHs, possibly due to using a different waveform model or the higher SNR of the GW150914-like injections.
@@
Fortunately, we do not require an accurate and universal “minimum detectable eccentricity” to establish the astrophysical implications of detecting eccentric NSBHs; we merely require an arbitrary eccentricity threshold above which we are confident that eccentricity measurement would be achieved. In accordance with standard practice for binary BH eccentricity measurability, we choose $e_{22,10}=0.05$ (Lower et al., 2018; Zevin et al., 2021). In the simulations of Stegmann and Klencki (2025), most of the NSBH mergers have eccentricities above this threshold: $\approx 80\%$ of NSBHs have $e_{22,10}\gtrsim 0.05$. Due to the low detectability of highly-eccentric inspiral-dominated systems, though, only $2.8\%$ of these are detected in our mock search. We therefore set the “above threshold” eccentric fraction, i.e. the fraction of systems that we are confident are detectable and measurably eccentric, to $\xi_{\mathrm{thresh}}=0.80\times 0.028=0.0224$. Using the same approach and notation as Zevin et al. (2021), in Figure 3, we plot the symmetric $95\%$ credible intervals around $\beta_{t}$, the branching fraction of field triples to the total population of observed NSBHs. If the number of observed NSBHs is $N_{\mathrm{obs}}=3$, and if GW200105 is truly eccentric so $N_{\mathrm{ecc}}=1$, then field triples may constitute $\geq 40\%$ of observed NSBHs.
@@
There are likely more than three NSBHs in the catalogue so far. In addition to GW190814, which may be an NSBH with an unusually massive NS secondary, GWTC-2 included the low-significance NSBH candidate GW190426_152155 (Abbott and others, 2021b). In GWTC-2.1, one confident candidate, GW190917_114630, has a secondary consistent with being an NS ($m_{2}\approx 2.1$ M), although it is identified by the search pipeline as a binary BH; there are two further low-significance NSBH candidates, GW190531_023648 and GW190917_114630 (Abbott et al., 2024). In GWTC-3, in addition to GW200115 and GW200105, two potential NSBHs—GW191219_163120 and GW200210_092254—were reported, although the former has an uncertain astrophysical probability, and the latter, like GW190814, has an ambiguous secondary mass (Abbott and others, 2023). In GWTC-4, both GW230529 and confident NSBH candidate GW230518_125908 were reported (The LIGO Scientific Collaboration et al., 2025a); however, the latter was not included in downstream analyses (e.g., the rates and populations analysis (The LIGO Scientific Collaboration et al., 2025b)) due to being detected outside of official observing, during the engineering run.
@@
We include in Figure 3 a vertical red band spanning $3\leq N_{\mathrm{obs}}\leq 11$, demonstrating the uncertainty on the number of NSBH detections. With $N_{\mathrm{obs}}=11$ and $N_{\mathrm{ecc}}=1$, the lower $90\%$ credible limit on $\beta_{t}$ is pushed down to $\approx 10\%$, while the detections remain consistent with $100\%$ of observed NSBHs originating in field triples at $90\%$ credibility. If no more eccentric NSBHs are detected, the upper $90\%$ credible limit on $\beta_{t}$ can begin to be constrained to <1<1 when we reach $N_{\mathrm{obs}}=157$.